# mycolorRGB219, 48, 122

### v1

@@ Abstract
We study phase separation between coexisting active and passive fluids in three-dimensions, using both numerical simulation and experiments. Chaotic flows of the active phase drive giant interfacial deformations and cause the co-existing phases to interpenetrate, generating a continuously reconfiguring bicontinuous steady-state morphology that persists over the lifetime of the active fluid. We demonstrate how activity controls the structure of the bicontinuous network. Quantitative analysis reveals the dominance of dynamical steady-state sheet-like interfaces, in marked difference from the transient bicontinuous structures observed in passive liquid-liquid phase separation, where saddle-like surfaces dominate. These results demonstrate how active stresses suppress the coarsening dynamics of conventional phase separation, generating steady-state reconfigurable morphologies not accessible with conventional surface-modifying agents or through quenching of transient phase separated structures.
@@ 1 Introduction
Bicontinuous structures consist of two distinct interconnected and interpenetrating phases, where each phase spans the entire subvolume. By maximizing the interfacial area while retaining connectivity, bicontinuity underpins vital mechanical and chemical functionalities in diverse natural and synthetic systems. An example from living matter is the endoplasmic that spans the cell interior. This bicontinuous structure performs protein synthesis, lipid production, and molecular transport on its interconnected membrane, while the bone trabecular lattice balances the mechanical strength of the mineralized struts with nutrient transport through vascularized marrow channels (?, ?, ?, ?). In synthetic systems, spinodally decomposed alloys derive exceptional strength and corrosion resistance from their interpenetrating nanoscale phases, while block copolymer gyroids form porous electrodes that optimize transport in batteries (?, ?, ?, ?, ?).
@@
A robust strategy for the fabrication of bicontinuous materials uses phase separation, a process in which a homogeneous fluid mixture spontaneously demixes into two distinct phases  (?). The phase separation dynamics can generate a bicontinuous structure  (?, ?). This structure, however, is inherently transient as the two phases coarsen over time reaching bulk phase separation  (?, ?). Thus, preserving bicontinuity requires kinetic arrest of the coarsening process. Established strategies include suppressing the fluidity of one of the two phases, either by replacing one of the two fluids with an elastomer network or by polymerizing one of the phase-separating fluids during coarsening (?, ?, ?, ?). The interface itself can also be stabilized by adsorbing a high density of colloids on the interface to form mechanically stable “bijels”  (?, ?, ?). Block copolymers offer an alternate method for generating bicontinous structures through chemistry, but the scale of the resulting structure is controlled by the size of the constituent molecules (?, ?, ?, ?, ?). Here we describe a fundamentally different pathway for generating robust and tunable bicontinuous structures that uses active fluids and relies on mechanical, rather than chemical, control of soft interfaces.
Active fluids are composed of energy-consuming molecular units that collectively generate autonomous chaotic flows  (?, ?, ?, ?, ?). Such flows and associated active stresses can power non-equilibrium dynamics and active assembly of passive colloids or fibers. They can also drive the deformation and actuation of soft interfaces and membranes  (?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?). We show how active flows continuously agitate the three-dimensional phase-separating mixture, transforming the conventional coarsening dynamics into a continuously reconfiguring bicontinuous network, which persists for the lifetime of the active fluid and is stable over a wide range of volume fractions. The bicontinuous structure is controlled by activity and surface tension. The interfacial curvature of active bicontinuous networks is dominated by sheet-like geometries, in contrast to passive phase-separating fluids dominated by minimal saddle-like surfaces  (?, ?, ?, ?, ?). The three-dimensional geometry studied here is qualitatively different from previous work which explored the role of microtubule-based activity on phase separation of PEO/dextran mixtures. One set of studies explored very thin chamber where the sample was effectively two dimensional  (?, ?). Another study explored the effect of activity on the shape and stability of a bulk-separated interface in a three dimensional sample  (?). Here, in contrast, we explore the morphology of the late stage phase separated state and demonstrate the stability of a remarkable bicontinuous structure.
@@
Figure 1: Active 3D phase separation yields a percolating network-like structure. (A-B) Evolution of the morphology of active phase separation in 3D computer simulations from the early state to the steady state. Red represents the active phase ($\phi>0.5$). Simulation parameters: activity $|\alpha|=1.30,$ surface tension $\gamma=1.0,$ and active fraction $\phi_{a}=0.40$. (C) Experimental morphology of the active phase (red) at steady state. The sample contains 92 nM KSA, 1.8% PEO, 1.8% dextran, and $\phi_{a}=0.38$, imaged at time $t=1.5$ h. (D) An $x$-$y$ cross-section of a phase-separating sample. The white arrows indicate the active flows in the dextran phase (red), obtained from Particle Imaging Velocimetry. (E) MTs (cyan) partition into the dextran phase (red). (F) MTs in the dextran phase form bundles due to the depletion effect. Kinesin clusters (KSA) bind to MT bundles driving their extension.
@@ 2.1 Theoretical model of active phase separation
We model an immiscible mixture of active and passive fluids in three-dimension within a continuum framework that couples Cahn-Hilliard phase separation to active nematodynamics. The concentration field $\phi(\mathbf{r})$ distinguishes the passive ($\phi=0$) and the active ($\phi=1$) phase, while the nematic tensor $\mathbf{Q}(\mathbf{r})$ captures the local orientational order within the active fluid. These fields are coupled through the incompressible Stokes flow $\mathbf{v}(\mathbf{r})$ and its spatial gradients (Methods 4.1)  (?, ?, ?).
@@
In passive systems, the well-mixed state is unstable over a range of parameters, leading to a spinodal instability with structures that coarsen over time. This reduces the interfacial energy, ultimately resulting in bulk phase separation  (?, ?). Activity profoundly alters this process  (?). At high activity, turbulent-like flows stir the fluid and impede the thermodynamically driven phase separation  (?, ?). Here we study intermediate activities that do not suppress phase separation, but arrest coarsening and influence the morphology of the phase separated state. In two dimensions, such moderate activity can generate percolating filamentary networks  (?). These networks have a dynamical steady-state, with a characteristic length scale controlled by the balance of activity and surface tension.
2.2 Active phase separation
We study the behavior of the above-described theoretical model in three dimensions. Starting from an initial uniform well-mixed state, the mixture quickly reaches a phase separated structure with a characteristic length scale. The steady-state has a network-like morphology that percolates throughout the sample (Fig. 1B). The emerging labyrinthine structure consists of junction points that are connected by thread-like liquid bridges. The network is not static. The threads connecting different junctions appear to continuously rupture, while new ones are generated such microscopic dynamics generates a “living” continuously reconfiguring network-like structure.
To experimentally test these predictions, we combined a passive phase-separating polymer mixture of polyethylene oxide (PEO) and dextran with an active fluid (Materials and Methods)  (?, ?, ?). The activity is generated by microtubules (MTs) and streptavidin-bound clusters of kinesin molecular motors (KSA) (Fig. 1F). Both protein-based components partition into the dextran-rich phase. The kinesin clusters bind to adjacent MTs, transforming the chemical energy from adenosine triphosphate (ATP) hydrolysis into interfilament sliding and bundle extension, which in turn drives large-scale turbulent flows.
@@ 2.2 Active phase separation
Figure 2: Bicontinuous morphology persists over a wide range of volume fractions. (A) Connected active (red) and passive phase (blue) in simulations, with $\phi_{a}=0.40$ (B) Connectedness factor $f_{c}$ of the active and passive phases as a function of the volume fraction of the active phase $\phi_{a}$. The range where $f_{c}\approx 1$ for both phases highlighted in pink indicates bicontinuity. Simulation parameters: $|\alpha|=0.70,\gamma=1.0$. (C) Morphology of the connected active (red) and passive phase (blue) in the experiment. $\phi_{a}=0.38$. (D) The connectedness fraction $f_{c}\approx 1$ for intermediate values of $\phi_{a}$, indicating bicontinuity. Samples initially contain 183 nM KSA, 1.8% PEO, and 1.8% dextran.
Figure 3: Active width $w_{a}$ decreases with activity and increases with surface tension. (A-B) $w_{a}$ decreases with activity $|\alpha|$ in simulations and with KSA concentration in experiments. (C-D) $w_{a}$ increases with surface tension $\gamma$ in simulations and with capillary length in experiments. In (A-D) the color of the dots represents the volume fraction $\phi_{a}$ of the active phase (see color bar) and shows that $w_{a}$ is essentially independent of $\phi_{a}$. (E-G) 3D visualization of the dextran phase (red) in experiments with increasing KSA concentration, showing a corresponding decrease of $w_{a}$. (H-J) 3D visualization of the dextran phase (red) in experiments with increasing capillary length, showing a corresponding increase of $w_{a}$. The experimental samples (B, E-G) contain initially 1.8% PEO and 1.8% dextran and varying KSA concentrations, 42 nM (E), 92 nM (F), and 183 nM (G). Samples (D, H-J) initially contain 183 nM KSA and varying PEO and dextran concentrations, both 1.8% (H), 1.9% (I), and 2.0% (J).
@@ 2.3 Characterizing the bicontinuous morphology
Numerical simulations and experiments demonstrate that 3D active phase separation generates a steady-state dynamical structure. Both active and passive phases form distinct yet intercalated system-spanning networks. Taken together, they generate a bicontinuous structure (Fig. 2 A,C). To analyze the bicontinuity we threshold the 3D structure into passive and active regions, corresponding to $\phi<0.5$ and $\phi>0.5$, respectively. The connectedness fraction, $f_{c}$, of each phase is defined as the volume of the largest connected active/passive component divided by the total volume of that component  (?).
@@
To determine the range of bicontinuity, we vary the fraction of the active phase $\phi_{a}$, while keeping the activity and surface tension fixed. Bicontinuous structures, defined as those where $f_{c}\approx 1$ for both phases, are observed from $0.3<\phi_{a}<0.6$ (Fig. 2 B). For $\phi_{a}<0.3$, bicontinuity is broken as the active domains disintegrate into isolated droplets. For $\phi_{a}>0.6$, the passive domains are broken up by active flows and become disconnected. The range of stable bicontinuous structures is controlled by activity and surface tension (Fig. S5). For example, doubling the activity generates stable bicontinuous structures over the entire regime of phase separation $0.0<\phi_{a}<1.0$.
@@
Motivated by these predictions, we have explored the phase space of stable bicontinuous networks in experiments. We use a phase recombination method to keep the polymer concentration (surface tension) and the motor protein concentration (activity) fixed within each phase, while the active fraction $\phi_{a}$ varies between 0.16-0.60. We (1) prepare an initial mixture with controlled concentrations of PEO, dextran and KSA, then (2) centrifuge the sample to generate bulk phase separation, and finally (3) recombine the two separated phases into a new mixture with chosen $\phi_{a}$ (Materials and Methods). The active (dextran) and passive (PEO) regions are differentiated using fluorescence intensity (Materials and Methods).
@@
In experiments, the bicontinous structure is observed over the range of volume fractions $0.3<\phi_{a}<0.5$ (Fig. 2D). For active fractions greater than $0.5$, the passive phase disintegrates into droplets, while for fractions less than $0.3$ the active phase breaks up into droplets, significantly decreasing $f_{c}$. Notably, the range of stable bicontinuous structures is not centered around $\phi_{a}=0.5$: bicontinuity persists for a low volume fraction of the active phase $\phi_{a}\approx 0.3$. In comparison, it disappears as soon as the active phase reaches the majority $\phi_{a}\approx 0.5$. This asymmetry suggests that the active phase containing MTs remains connected at a very low volume fraction, while the passive PEO-rich/MT-devoid phase more easily breaks apart into the droplets. Intriguingly, similar asymmetry was observed in the directional invagination and break-up of a bulk three-dimensional active interface, where it was attributed to the nematic elasticity of the active fluid  (?).
@@ 2.4 Controlling the morphology: active width
Simulations show that the morphology of the bicontinuous network is controlled by activity and surface tension. To characterize the structure, we measure the width $w_{a}$ of the active network, defined as the surface to volume ratio of the largest connected active domain. We find that $w_{a}$ increases with decreasing activity and with increasing equilibrium surface tension. Furthermore, $w_{a}$ is essentially independent of the volume fraction $\phi_{a}$ (Fig. 3A, C), suggesting that intrinsic dynamical active processes control the length scale of the bicontinuous morphology.
To experimentally test these predictions, we vary the KSA concentration, which is a proxy for activity, and the PEO/dextran concentration, which controls the surface tension. Due to a density difference, the dextran phase gradually sediments due to gravity over tens of minutes, causing a variation in $w_{a}$ (Fig. S3). To minimize this effect, we continuously invert the top and bottom of the chamber by slowly rotating the samples. The sample is then moved from the rotation stage and imaged for a maximum duration of 5 minutes (Materials and Methods). The effective surface tension is estimated from measurements of the capillary length for different PEO/dextran concentrations (Fig. S1). The experimental trends follow the predictions of the numerical model: increasing surface tension increases $w_{a}$; increasing activity decreases $w_{a}$ (Fig. 3 A-D). Furthermore, $w_{a}$ is largely independent of the volume fraction.
The behavior of the width $w_{a}$ suggests that the morphology is controlled by two competing mechanisms: the stretching of the interface due to extensile active flows and its relaxation controlled by surface tension. Extensile active forces are known to align the nematic director with the interface (?, ?, ?). This alignment then drives active shear flows that tangentially stretch the interface. The stretching is countered by surface tension that tends to shorten the interface. In general we expect activity to also modify the surface tension, as shown for instance in scalar models of active phase separation  (?). Here we assume that the main effect of activity is to build up local nematic order and align the liquid crystal with the interface, and use the bare (passive) surface tension in the scaling arguments below. In the isotropic liquid crystalline phase, the active stretching rate can be approximated as
@@
where $\eta$ is the fluid viscosity, $S$ is the scalar nematic order parameter and $\alpha_{c}$ is the threshold above which activity drives local nematic order (?, ?) (see Methods Section). We approximate the scalar order parameter with the mean value generated by activity, $S\sim\sqrt{(|\alpha|-\alpha_{c})}$ (?, ?). In three-dimensions, the rate at which the interface relaxes due to surface tension is given by
@@
Figure 4: Steady state width $w_{a}=\text{Volume}/\text{Surface Area}$ of the connected active domain in the bicontinuous region as a function of activity $|\alpha|$ and equilibrium surface tension $\gamma$, for a fixed active fraction $\phi_{a}=0.40$. The main figure shows that the data collapse when plotted versus $(|\alpha|-\alpha_{c})/\gamma^{2/3}$ (Eq. 4). The black dashed line shows the best fit (via linear regression) to $w_{a}\sim(|\alpha|-\alpha_{c})/\gamma^{2/3})^{\nu}$ for $\nu=-1.36\pm 0.04$. The inset shows the unscaled active width as a function of activity, for various values of $\gamma$. Fixed parameters: $K=24.0,\alpha_{c}=2\eta r/\lambda\Gamma=0.2$.

### v2

@@ Abstract
We study phase separation between coexisting active and passive fluids in three-dimensions, using numerical simulation and experiments. Chaotic flows of the active phase drive giant interfacial deformations, causing the co-existing phases to interpenetrate and generate a continuously reconfiguring bicontinuous morphology which persists over the lifetime of the active fluid. Active bicontinuous structures are dominated by sheet-like interfaces, in marked difference from passive liquid-liquid phase separation which is controlled by saddle-like surfaces. Activity and surface tension control the length scale of the bicontinuous structure. These results demonstrate how active stresses suppress the coarsening of conventional phase separation, generating steady-state reconfigurable morphologies not accessible with conventional surface-modifying agents or through quenching of transient phase separated structures.
@@ 1 Introduction
Bicontinuous structures consist of two distinct interconnected and interpenetrating phases, where each phase spans the entire sample. By maximizing the interfacial area while retaining connectivity, bicontinuity underpins vital mechanical and chemical functionalities in diverse natural and synthetic systems. For example, in living matter, the interconnected endoplasmic reticulum that spans the cell interior performs protein synthesis, lipid production, and molecular transport  (?, ?). The bone trabecular lattice balances the mechanical strength of the mineralized struts with nutrient transport through vascularized marrow channels  (?, ?). In synthetic systems, spinodally decomposed alloys derive exceptional strength from their interpenetrating nanoscale phases, while block copolymer gyroids form porous electrodes that optimize transport in batteries (?, ?, ?, ?).
@@
A homogeneous fluid mixture spontaneously demixes into two distinct phases  (?), a dynamical process that can generate a bicontinuous structure  (?, ?). In systems in thermal equilibrium, however, bicontinuous structure is inherently transient as the two phases coarsen over time reaching bulk phase separation  (?, ?). Thus, preserving bicontinuity requires kinetic arrest of the coarsening. Established strategies include suppressing the fluidity of one of the two phases. This can be achieved by replacing one of the two fluids with an elastomer network (?), by polymerizing one of the phase-separating fluids during coarsening (?, ?, ?), or by vitrifying it through rapid cooling (?). Alternatively, the interface itself can be stabilized by adsorbing a high density of colloids on the interface to form mechanically stable “bijels”  (?, ?, ?). Finally, block copolymers offer an alternate method for generating bicontinous structures through chemistry, but the scale of the resulting structure is controlled by the size of the constituent molecules  (?, ?, ?).
@@
Figure 1: Active 3D phase separation yields a percolating network-like structure. (A-B) Morphology of active phase separation in computer simulations from the early state to the steady state. Red represents the active phase ($\phi>0.5$). Simulation parameters: activity $|\alpha|=1.30,$ surface tension $\gamma=1.0,$ and active fraction $\phi_{a}=0.40$. (C) Experimental morphology of the active phase (red) at steady state. (D) An $x$-$y$ cross-section of a phase-separating experimental sample. The white arrows indicate the active flows in the dextran phase (red), obtained from Particle Imaging Velocimetry. (E) MTs (cyan) partition into the dextran phase (red). (F) MTs in the dextran phase form bundles due to the depletion effect. Kinesin clusters (KSA) bind to MT bundles driving their extension. The experimental samples (C-E) contain 1.8% PEO and 1.8% dextran, $\phi_{a}=0.37$, and are imaged at time $t=1.5$ h. The KSA concentration is 183 nM (C) and 92 nM (D, E).
Here we describe a fundamentally different pathway for generating robust and tunable bicontinuous structures that uses active fluids and relies on mechanical, rather than chemical, control of soft interfaces. Active fluids are composed of energy-consuming molecular units that collectively generate autonomous chaotic flows  (?, ?, ?, ?, ?). Such flows and associated active stresses can power non-equilibrium dynamics and active assembly of passive colloids or fibers (?, ?). They can also drive the deformation and actuation of soft interfaces and membranes  (?, ?, ?, ?, ?, ?, ?, ?, ?, ?). Here we show how active flows can transform conventional coarsening dynamics into a continuously reconfiguring bicontinuous active network, which persists for the lifetime of the active fluid and is stable over a wide range of volume fractions. The bicontinuous structure is controlled by activity and surface tension. The interfacial curvature of the active bicontinuous network is dominated by sheet-like geometries, in contrast to passive phase-separating fluids where bicontinuous structures are dominated by minimal saddle-like surfaces  (?, ?, ?, ?, ?).
The three-dimensional geometry studied here is qualitatively different from previous work which explored the role of microtubule-based activity on phase separation of PEO/dextran mixtures. One class of studies explored active phase separation in very thin chambers where the sample was effectively two dimensional  (?, ?). Another study examined the effect of activity on the shape and stability of a bulk-separated interface in a three dimensional sample  (?). Here, in contrast, we explore the morphology of the late stage phase separated state and demonstrate the stability of bicontinuous structures.
@@ 2.1 Theoretical model of active phase separation
We model an immiscible mixture of active and passive fluids in three dimensions within a continuum framework that couples Cahn-Hilliard phase separation to active nematodynamics. The concentration field $\phi(\mathbf{r})$ distinguishes the passive ($\phi=0$) and the active ($\phi=1$) phase, while the nematic tensor $\mathbf{Q}(\mathbf{r})$ captures the local orientational order within the active fluid. These fields are coupled through the incompressible Stokes flow $\mathbf{v}(\mathbf{r})$ and its spatial gradients (Methods 4.1)  (?, ?, ?).
@@
In passive systems, the well-mixed state is unstable over a range of parameters, leading to a spinodal instability that generates spatially modulated structures that coarsen over time. Such dynamics reduces the interfacial energy, ultimately resulting in bulk phase separation  (?, ?). Activity profoundly alters this process  (?). At high activity, turbulent-like flows stir the fluid and impede the thermodynamically driven phase separation  (?, ?). Intermediate activities do not suppress phase separation, but arrest coarsening and influence the morphology of the phase separated state. In two dimensions, moderate activity can generate percolating dynamical filamentary networks with a characteristic length scale controlled by the balance of activity and surface tension  (?).
We study the active-passive mixture at intermediate activities, using the numerical model. Starting with an initially uniform state, we observe the emergence of spatially modulated structures, which reach a length scale that plateaus over time (Fig. 1A, B, Movie S1). The steady-state structures have a labyrinth like appearance, consisting of junction points that are connected by thread-like liquid bridges. The network is not static. The threads connecting different junctions appear to continuously rupture while new ones are generated. Such microscopic dynamics generates a “living” continuously reconfiguring network-like structure.
2.2 Experimental realization of active phase separation
To experimentally study active-passive phase separation, we combined a previously described passive phase-separating polymer mixture of polyethylene oxide (PEO) and dextran with an active fluid (Methods 4.2)  (?, ?, ?). The activity is generated by microtubules (MTs) and streptavidin-bound clusters of kinesin molecular motors (KSA) (Fig. 1F). Both protein-based components partition into the dextran-rich phase (Fig. 1E). The kinesin clusters bind to adjacent MTs, transforming the chemical energy from adenosine triphosphate (ATP) hydrolysis into interfilament sliding and bundle extension, which in turn drives large-scale turbulent flows (Fig. 1D).
@@ 2.2 Experimental realization of active phase separation
Figure 2: Bicontinuous morphology persists over a wide range of volume fractions. (A) Connected active (red) and passive phase (blue) in simulations, with $\phi_{a}=0.40$ (B) Connectedness factor $f_{c}$ of the active and passive phases as a function of the volume fraction of the active phase $\phi_{a}$. The range where $f_{c}\approx 1$ for both phases highlighted in pink indicates bicontinuity. Simulation parameters: $|\alpha|=0.70,\gamma=1.0$. (C) Morphology of the connected active (red) and passive phase (blue) in the experiment. $\phi_{a}=0.37$. (D) The connectedness fraction $f_{c}\approx 1$ for intermediate values of $\phi_{a}$, indicating bicontinuity. Experimental samples initially contain 183 nM KSA, 1.8% PEO, and 1.8% dextran.
Figure 3: Active width $w_{a}$ decreases with activity and increases with surface tension. (A-B) $w_{a}$ decreases with activity $|\alpha|$ in simulations and with KSA concentration in experiments. (C-D) $w_{a}$ increases with surface tension $\gamma$ in simulations and with capillary length in experiments. In (A-D) the color of the dots represents the volume fraction $\phi_{a}$ of the active phase (see color bar) and shows that $w_{a}$ is essentially independent of $\phi_{a}$. (E-G) 3D visualization of the dextran phase (red) in experiments with increasing KSA concentration, showing a corresponding decrease of $w_{a}$. (H-J) 3D visualization of the dextran phase (red) in experiments with increasing capillary length, showing a corresponding increase of $w_{a}$. The experimental samples (B, E-G) contain 1.8% PEO and 1.8% dextran. KSA concentrations are increased from 42 nM (E), 92 nM (F), to 183 nM (G). Samples (D, H-J) contain 183 nM KSA. PEO and dextran concentrations are both increased from 1.8% (H), 1.9% (I), to 2.0% (J).
@@ 2.3 Characterizing the bicontinuous morphology
Numerical simulations and experiments demonstrate that 3D active phase separation generates a steady-state dynamical structure. Active and passive phases form distinct yet intercalated system-spanning networks. Taken together, they generate a bicontinuous structure (Fig. 2 A,C). To analyze the bicontinuity in simulation, we threshold the 3D structure into passive and active regions, corresponding to $\phi<0.5$ and $\phi>0.5$, respectively. The connectedness fraction, $f_{c}$, of each phase is defined as the volume of the largest connected active/passive component divided by the total volume of that component  (?).
@@
To determine the range of bicontinuity, we vary the fraction of the active phase $\phi_{a}$, while keeping the activity and surface tension fixed. Bicontinuous structures, defined as those where $f_{c}\approx 1$ for both phases, are observed from $0.3<\phi_{a}<0.6$ (Fig. 2 B) in simulation. For $\phi_{a}<0.3$, bicontinuity is broken as the active domains disintegrate into isolated droplets. For $\phi_{a}>0.6$, the passive domains are broken up by active flows and become disconnected. The range of stable bicontinuous structures is controlled by activity and surface tension (Fig. S5). For example, doubling the activity generates stable bicontinuous structures over the entire regime of phase separation $0.0<\phi_{a}<1.0$.
@@
Motivated by these predictions, we have explored the phase space of stable bicontinuous networks in experiments. We use a phase recombination method to keep the polymer concentration (surface tension) and the motor protein concentration (activity) fixed within each phase, while the active fraction $\phi_{a}$ varies between 0.16-0.60. We (1) prepare an initial mixture with controlled concentrations of PEO, dextran and KSA, then (2) centrifuge the sample to generate bulk phase separation, and (3) recombine the two separated phases into a new mixture with chosen $\phi_{a}$ (Methods 4.2).
@@
In experiments, the bicontinous structure for the particular is observed over the range of volume fractions $0.3<\phi_{a}<0.5$ (Fig. 2D). For active fractions greater than $0.5$, the passive phase disintegrates into droplets, while for fractions less than $0.3$ the active phase breaks up into droplets, significantly decreasing $f_{c}$. Notably, the range of stable bicontinuous structures is not centered around $\phi_{a}=0.5$: bicontinuity persists for a low volume fraction of the active phase $\phi_{a}\approx 0.3$. In comparison, it disappears as soon as the active phase reaches the majority $\phi_{a}\approx 0.5$. This asymmetry suggests that the active phase containing MTs remains connected at a very low volume fraction, while the passive PEO-rich/MT-devoid phase more easily breaks apart into the droplets. This asymmetry is more pronounced in experiments than simulations. Intriguingly, similar asymmetry was observed in the directional invagination and break-up of a bulk three-dimensional active interface, where it was attributed to the nematic elasticity of the active fluid  (?).
@@ 2.4 Controlling the morphology: active width
Simulations show that the morphology of the bicontinuous network is controlled by activity and surface tension. To characterize this structure, we measure the width $w_{a}$ of the active network, defined as the volume to surface ratio of the largest connected active domain. We find that $w_{a}$ increases with decreasing activity and with increasing equilibrium surface tension. Furthermore, $w_{a}$ is essentially independent of the volume fraction $\phi_{a}$ (Fig. 3A, C), suggesting that intrinsic active processes control the length scale of the bicontinuous morphology.
To experimentally test these predictions, we vary the KSA concentration, which is a proxy for activity, and the PEO/dextran concentration, which controls the surface tension. Due to a density difference, the dextran phase gradually sediments due to gravity over tens of minutes, causing a variation in $w_{a}$ (Fig. S3). To minimize this effect, we continuously invert the top and bottom of the chamber by slowly rotating the samples. The sample is then moved from the rotation stage to the microscope and imaged for a maximum duration of 5 minutes (Methods 4.2). The effective surface tension is estimated from measurements of the capillary length for different PEO/dextran concentrations (Fig. S1). The experimental trends follow the predictions of the numerical model: increasing surface tension increases $w_{a}$; increasing activity decreases $w_{a}$ (Fig. 3 A-D). Furthermore, $w_{a}$ is largely independent of the volume fraction.
The behavior of the width $w_{a}$ suggests that the morphology is controlled by two competing mechanisms: the stretching of the interface due to extensile active flows and its relaxation controlled by surface tension. Extensile active forces are known to align the nematic director with the interface  (?, ?, ?). This alignment then drives active shear flows that tangentially stretch the interface. The stretching is countered by surface tension that favors minimizing the interface. In general we expect activity to also modify the surface tension, as shown for instance in scalar models of active phase separation  (?). Here we assume that the main effect of activity is to build up local nematic order and align the liquid crystal with the interface, and use the bare (passive) surface tension in the scaling arguments below. In the isotropic liquid crystalline phase, the active stretching rate can be approximated as
@@
where $\eta$ is the fluid viscosity, $S$ is the scalar nematic order parameter and $\alpha_{c}$ is the threshold above which activity drives local nematic order (?, ?) (Methods 4.1). We approximate the scalar order parameter with the mean value generated by activity, $S\sim\sqrt{(|\alpha|-\alpha_{c})}$ (?, ?). In three-dimensions, the rate at which the interface relaxes due to surface tension is given by
@@
Figure 4: Scaling law for the width $w_{a}$ in simulations. Steady state width $w_{a}=\text{Volume}/\text{Surface Area}$ of the connected active domain in the bicontinuous region as a function of activity $|\alpha|$ and equilibrium surface tension $\gamma$, for a fixed active fraction $\phi_{a}=0.40$. The main figure shows that the data collapse when plotted versus $(|\alpha|-\alpha_{c})/\gamma^{2/3}$ (Eq. 4). The black dashed line shows the best fit (via linear regression) to $w_{a}\sim(|\alpha|-\alpha_{c})/\gamma^{2/3})^{\nu}$ for $\nu=-1.36\pm 0.04$. The inset shows the unscaled active width as a function of activity, for various values of $\gamma$. Fixed parameters: $K=24.0,\alpha_{c}=2\eta r/\lambda\Gamma=0.2$.